Find a function h(t) such that h'(t) = -32t + v[0] for all t, h(0) = 20 and h(1) = 0. (Here v[0] is a constant to be determined.)
Please explain how to solve
Thank you

Answer by stanbon(75887)

Find a function h(t) such that h'(t) = -32t + v[0] for all t, h(0) = 20 and h(1) = 0. (Here v[0] is a constant to be determined.)
-----
Integrate h'(t) to get:
h(t) = -16t^2 + v(0)t + C
---
h(0) = -16*0^2+v(0)*0 + C = 20
So, C = 20
------------
h(1) = -16(1^2) + v(0)*1 + 20 = 0
-16 + v(0) + 20 = 0
v(0) = -4
------
Ans: h(t) = -16t^2 -45 + 20
================
Cheers,
Stan H.
--------------